The length of a rectangle is 6 ft longer than its width. The perimeter is 84 ft . What is the length?
84=2w+2l then I forget how to do this. Thank you.

Answer by Earlsdon(6294)

That's a good start:
2W + 2L = 84 But...L = W+6, so substitute this for L.
2W + 2(W+6) = 84 Simplify and solve for W.
2W + 2W + 12 = 84 Combine like-terms.
4W + 12 = 84 Subtract 12 from both sides.
4W = 72 Divide both sides by 4.
W = 18
The width is 18 ft.
The length, L = W+6 = 18+6 = 24 ft.